### v1

# Two-stage Generative Question Answering on Temporal Knowledge Graph Using Large Language Models

Abstract
Temporal knowledge graph question answering (TKGQA) poses a significant challenge task, due to the temporal constraints hidden in questions and the answers sought from dynamic structured knowledge. Although large language models (LLMs) have made considerable progress in their reasoning ability over structured data, their application to the TKGQA task is a relatively unexplored area. This paper first proposes a novel generative temporal knowledge graph question answering framework, GenTKGQA, which guides LLMs to answer temporal questions through two phases: Subgraph Retrieval and Answer Generation. First, we exploit LLM’s intrinsic knowledge to mine temporal constraints and structural links in the questions without extra training, thus narrowing down the subgraph search space in both temporal and structural dimensions. Next, we design virtual knowledge indicators to fuse the graph neural network signals of the subgraph and the text representations of the LLM in a non-shallow way, which helps the open-source LLM deeply understand the temporal order and structural dependencies among the retrieved facts through instruction tuning. Experimental results demonstrate that our model outperforms state-of-the-art baselines, even achieving 100% on the metrics for the simple question type.
Two-stage Generative Question Answering on Temporal Knowledge
Graph Using Large Language Models
[…]
Figure 1: Examples of the responses of LLM and GenTKGQA to the simple and complex temporal questions.
1 Introduction
Real-world knowledge is frequently updated rather than static Erxleben et al. (2014); Boschee et al. (2015), e.g., (Obama, hold_position, President) is merely valid only for a certain period [2009, 2016]. Hence, the temporal knowledge graph (TKG) is proposed as a database for storing dynamic structured facts associated with timestamps, denoted as (subject, relation, object, timestamp). Temporal knowledge graph question answering (TKGQA) aims to answer a natural question with explicit or implicit temporal constraints based on the TKG, e.g., "Who held the position of president (in 2017) or (after Obama)?". Due to the temporal constraints hidden in questions and the answers being sought from dynamic structured knowledge, TKGQA is one of the most challenging QA tasks.
Recently, large language models (LLMs) have shown strong competitiveness in various fields Fei et al. (2023); Ye et al. (2023). Some researchers explore the reasoning capability of LLMs for structured knowledge based on KGQA tasks Baek et al. (2023); Kim et al. (2023), and some works examine the temporal reasoning capabilities of LLMs through time-sensitive QA tasks Chen et al. (2021a); Tan et al. (2023). Intuitively, LLMs have the ability to deal with temporal structured knowledge. Based on the above findings, we attempt to utilize LLMs for the TKGQA task and summarize the following two challenges: 1) Question-relevant Subgraph Retrieval. A common practice to enhance the LLM’s domain-specific reasoning capability is to input query-relevant information as additional knowledge into the LLMs Hu et al. (2023). As shown in Figure 1, ChatGPT OpenAI (2023) cannot answer temporal questions directly, but it can answer simple questions when the relevant context is provided. However, finding facts relevant to the problem is a struggle due to the large search space with both structural and temporal dimensions. For example, for the complex temporal question "Who held the position of president after Obama?", the structural link hold position between the entities and the temporal constraints after 2016 in the problem are unknown, so directly finding all relevant facts about the given entities Obama and President is bound to introduce too much noisy information. How to accurately retrieve relevant facts from a two-dimensional space is the first challenge. 2) Complex-type Question Reasoning. Recent research about LLM-based KG reasoning mostly inputs structured knowledge in the natural text form into the task prompt and reasons about the answers in a training-free method Yang et al. (2024). However, this approach fuses subgraph information with the LLM in a shallow manner, which limits the inference performance on the complex question type. As illustrated in Figure 1, ChatGPT is unable to understand the chronological order of the relevant facts and answers incorrectly on the complex question type "after". How subgraph information can be integrated into LLM representations in a non-superficial way to simulate structured reasoning remains an open question.
Hence, we propose GenTKGQA, a novel generative temporal knowledge graph question answering framework consisting of two phases, subgraph retrieval and answer generation, which is used to address the above two challenges, respectively. At the first phase, we find that the structural and temporal scope of the subgraph is determined by the relation links and the temporal constraints in the question, respectively. Therefore, we use a divide-and-conquer strategy to reduce the subgraph search space by decomposing the complex subgraph retrieval problem into two subtasks, namely, relation ranking and time mining. Then, we utilize the LLM’s internal knowledge to mine structural connections between entities and time constraints in the problem without extra training. We only need to input few-shot examples into the prompt to accomplish subgraph retrieval of the entire data. At the second phase, we fine-tune the open-source LLM with instruction tuning to incorporate structural and temporal information of the subgraph in a non-shallow way. Recent works illustrate that fusing graph neural network (GNN) representations and language text representations can enhance the ability of LMs to perceive graph structure Zhang et al. (2022b). Thus, we design three novel virtual knowledge indicators to bridge the links between pre-trained GNN signals of the temporal subgraph and text representations of the LLMs, which guides the LLMs in deeply understanding the graph structure and improves their reasoning ability for complex temporal questions. Overall, our contribution can be summarized in the following four points:
1) We present a novel two-stage generative framework for the TKGQA task, which explores LLM’s temporal reasoning capabilities in the context of dynamic structured knowledge.
2) We motivate the LLM’s intrinsic knowledge to mine the temporal constraints and structural connections in the questions without extra training, which reduces the subgraph search space from both structure and time dimensions.
3) We design virtual knowledge indicators to fuse the GNN signals and text representations in a non-shallow way, which helps the open-source LLMs improve their reasoning on the complex question type through instruction tuning.
4) Experiment results show that GenTKGQA as a generative QA model performs consistently better than embedding-based QA methods, even reaching 100% on all metrics for the simple question type.
2 Related Work
2.1 TKGQA Methods
Temporal knowledge graph question answering (TKGQA) task aims to answer complex questions in the natural language format using entities and timestamps from the given TKG Jia et al. (2018, 2021); Saxena et al. (2021); Chen et al. (2023). Existing mainstream methods employ TKG embeddings to represent the entities, relations and timestamps, and use the TKGE scoring function to select the entity or time with the highest relevance as the answer Saxena et al. (2021). However, single embedding methods have difficulty handling complex reasoning problems with implicit time constraints. Therefore, recent methods try to incorporate other modules to improve the model performance on complex problems. Specifically, TSQA Chen et al. (2021b) proposes a contrastive approach to enhance the time sensitivity of the model. TempoQR Mavromatis et al. (2022) designs three modules, namely context, entity, and time-aware information, to enhance the incorporation of the TKG into questions. LGQA Liu et al. (2023b) applies a multi-hop message passing GNN layer to combine the global and local information of the given problem. Despite the effectiveness of these approaches, few works explore how approaches other than the embedded-based method can address the TKGQA task.
2.2 LMs for Temporal Question Answering
Language models (LMs) have exhibited strong performance on the question answering task Raffel et al. (2020). In recent years, some researchers have explored the temporal reasoning capabilities of LMs and propose several typical time-sensitive QA datasets. They focus on temporal question answering either within a closed-book setting to assess models’ internal memorization of temporal facts Liska et al. (2022); Dhingra et al. (2022), or within an open-book setting to evaluate models’ temporal understanding and reasoning capability over unstructured texts Zhang and Choi (2021); Chen et al. (2021a); Tan et al. (2023). In the context of the latter setting, some works propose to use the graph structure extracted from text to assist the model in determining the temporal order between events Mathur et al. (2022); Su et al. (2023); Yang et al. (2023), which is similar but fundamentally different from our work. These approaches aim to answer temporal questions based on the known natural text context. In contrast, our model focuses on structured temporal knowledge as auxiliary information that needs to be retrieved by the model.
2.3 LMs for KG Question Answering
How to combine LMs and KG for question answering has become a hot issue. Some works attempt to enhance question representation and relation matching with PLMs in the multi-hop KGQA task Saxena et al. (2020); Zhang et al. (2022a); Jiang et al. (2023b), but there is no interaction between the LM and KG representations. Other works try to use one modality to ground the other, i.e., using the encoded representation of a linked KG to augment the text representation Lin et al. (2019); Yang et al. (2019), or using the text representation of the PLM to enhance the graph reasoning model Feng et al. (2020a). The most recent approaches enable deep integration of the two modalities by jointly updating the GNN and LM representation Yasunaga et al. (2021); Zhang et al. (2022b).
However, the emergence of large language models (LLMs) has changed how LMs handle the KGQA task, which is divided into two main approaches: training-free and fine-tuning Yang et al. (2024). Recent works attempt to append query-relevant facts as the input prompt for LLMs and make inferences without extra training Baek et al. (2023); Jiang et al. (2023a); Kim et al. (2023); Li et al. (2024). Fine-tuning the full parameters of the LLM can be cost-prohibitive. Hence, KPE Zhao et al. (2023) enables knowledge integration by freezing PLM parameters and introducing trainable parameter adapters. ChatKBQA Luo et al. (2023) employs the LoRA Hu et al. (2022) technique to fine-tune open-source LLMs, achieving the logical query form generation. Applying LLMs to the temporal KGQA task remains an unexplored area.
3 Preliminaries
TKGQA.
A temporal knowledge graph (TKG) $\mathcal{G}:=(\mathcal{E},\mathcal{R},\mathcal{T},\mathcal{F})$ is a multi-relational, directed graph with timestamped edges between entities, where $\mathcal{E}$, $\mathcal{R}$ and $\mathcal{T}$ represent the sets of entities, relationships and timestamps, respectively. Each fact in the $\mathcal{G}$ can be represented as a quadruple $(s,r,o,t)\in\mathcal{F}$, corresponding to entity $s/o\in\mathcal{E}$, relation type $r\in\mathcal{R}$ and timestamp $t\in\mathcal{T}$. Given a natural language question $q$, TKGQA aims to extract an entity $s/o$ or a timestamp $t$ that correctly answers the question.
ICL and IT.
Applying LLM to the TKGQA task, the goal is to generate the answer $\mathcal{A}$ based on the input text sequence $\mathcal{S}$ and the LLM $\mathcal{M}$. $\mathcal{S}$ consists of several parts: the instruction prompt $\mathcal{I}$, the task-specific input prompt $\mathcal{Q}$, and the auxiliary demonstration prompt $\mathcal{D}$. In-context Learning (ICL) method is an efficient approach to employ LLMs to solve downstream tasks without extra training, the input sequence of ICL can be denoted as $\mathcal{S}=\mathcal{I}:\mathcal{D}:\mathcal{Q}$, where :::: means to concatenate the different prompts. Meanwhile instruction tuning (IT) aims to fine-tune LLMs to follow human instructions and accomplish the distinct tasks in the instruction prompt, the input sequence of IT can be denoted as $\mathcal{S}=\mathcal{I}:\mathcal{Q}:\mathcal{A}$.
4 Method
We apply the LLMs processing TKGQA task in a two-phase process, i.e., an ICL-based subgraph retrieval phase and an IT-based answer generation phase. At the first stage, we utilize internal knowledge of LLM for unlabeled subgraph retrieval. At the second stage, we incorporate external knowledge for structure-aware temporal inference.
Figure 2: The overall architecture of our proposed GenTKGQA can be divided into two stages, subgraph retrieval and answer generation. Given a temporal problem, we mark the entities provided, the implied time constraints and the links between the entities with underline, blue font and orange font, respectively.
4.1 Subgraph Retrieval
We split the complex subgraph retrieval problem into a relation ranking sub-task in the structural dimension and a time mining sub-task in the temporal dimension. With such a strategy, we only need to provide a small number of samples to complete the subgraph retrieval.
4.1.1 Relation Ranking
We aim to determine the structural scope of the subgraph $\mathcal{G}_{sub,i}$ by retrieving the corresponding relations from the candidate relation set $\mathcal{R}_{i}$ for each question $q_{i}$. Recent work has shown that LLMs can better handle the information extraction task as re-ranking agents Sun et al. (2023). Therefore, we feed the question $q_{i}$ and the candidate set $\mathcal{R}_{i}$ to the LLM to obtain the top $k$ relations $\mathcal{R}_{i,k}$ relevant to the question. Relations in $\mathcal{R}_{i}$ are linearized, i.e., [employer, member_of_sports_team, …, ]), and the retrievd relations can bridge the entities identified within the questions. The specific relation ranking prompt is shown in Appendix C.
4.1.2 Time Mining
We find that natural questions contain temporal constraints, either explicit or implicit, and we can easily determine the range of relevant facts in the temporal dimension by using explicit temporal constraints such as "in 2008", "at the year of 2012", etc. How to capture implicit temporal constraints is the key to improving the efficiency of searching for relevant facts. For example, for the complex temporal question "Who held the position of president after Obama?", the implicit temporal restriction is known to be (after 2016) based on the temporal validity (2009, 2016) of the fact (Obama, hold_position, President). We design specific prompt templates based on different answer types as well as consider the temporal validity of the question-containing facts (given entities and relations matched in the previous section) to get the temporal constraints. The details of the templates are shown in Appendix C.
Through the above process, we narrow down the search space of subgraphs and use relevant facts under structural and temporal constraints as additional knowledge to assist the LLM inference.
4.2 Answer Generation
In this section, we will discuss how to incorporate the knowledge retrieved in the previous section 4.1 into the LLM. Previous fundamental approaches to incorporate KG structural information focus on adding the knowledge to the input prompt in the text form, e.g., (subject, relation, object). However, incorporating query-relevant facts into LLMs in text form is not a good choice. The reason is that this shallow interaction does not enable the model to understand the structural dependencies and temporal order between facts, leading to weak temporal reasoning in the complex problem type.
Recent works show that language models can enhance their ability to perceive graph structures by incorporating knowledge representations expressed by graph neural networks (GNN) Zhang et al. (2022b). Inspired by this, we first extract the structural and temporal information of entities and relations with pre-trained temporal GNN embeddings. Then, we bridge the links between GNN and text representations through pre-designed virtual knowledge indicators. At last, we fine-tune the open-source LLM to deeply understand the temporal order and structural dependencies of the retrieved query-relevant facts.
4.2.1 Temporal GNN
Given the retrieved temporal subgraph $\mathcal{G}_{sub,z}$ of question $q_{z}$, we first initialise entity, relation and time representations in $\mathcal{G}_{sub,z}$ using the TKG embedding method Lacroix et al. (2020). Then, to fully explore the structural information among entities and relations of the temporal subgraph, we propose a temporal graph neural network (T-GNN), which is a variant of graph attention networks Velickovic et al. (2018). The important distinction between them is that T-GNN captures the correlation scores of neighbouring nodes by incorporating temporal embeddings. Therefore, T-GNN computes message $\mathbf{m}_{ij}$ between entities $e_{i}$ and $e_{j}$ as follows:
$\mathbf{m}_{ij}=\mathbf{W}_{m}(\mathbf{e}_{i}^{(l-1)}+\mathbf{r}_{ij}+\mathbf{t}_{ij}),$ (1)
where $\mathbf{e}_{i}^{(l-1)}$ is the entity representation of $e_{i}$ at the $l$-1 layer, $\mathbf{r}_{ij}$ and $\mathbf{t}_{ij}$ are the embeddings of the relation and timestamp connecting $e_{i}$ and $e_{j}$. $\mathbf{W}_{m}$ is a linear transformation. Next, the node representation $\mathbf{e}_{j}^{(l)}$ is calculated via message passing between neighbors on the $\mathcal{G}_{sub}$:
$\mathbf{e}_{j}^{(l)}=\sum_{i\in\mathcal{N}_{j}}\alpha_{ij}\mathbf{m}_{ij},$ (2)
here $\mathcal{N}_{j}$ represents the neighbor entities of the arbitrary node $e_{j}$, and $\alpha_{ij}$ denotes the attention values with $\mathbf{e}_{i}^{(l-1)}$ as query and $\mathbf{m}_{ij}$ as key:
$\alpha_{ij}=\frac{\exp\left(u_{ij}\right)}{\sum_{w\in\mathcal{N}_{j}}\exp\left(u_{wj}\right)},$ (3)
$u_{ij}=f_{n}((\mathbf{W}_{q}\mathbf{e}_{i}^{(l-1)})^{\top}\left(\mathbf{W}_{k}\mathbf{m}_{ij}\right)),$ (4)
$\mathbf{W}_{q},\mathbf{W}_{k}$ are linear transformations, $f_{n}$ is the RELU activation function. Through the above process, we obtain entity representations $\mathbf{e}_{j}$ with subgraph structural and temporal information. Following embedding-based TKGQA methods, we use the link prediction task to pre-train the graph neural network representations. Specifically, for each fact $(s,r,o,t)$ in the TKG, we generate a query $(s,r,[mask],t)$ or $([mask],r,o,t)$ by masking the object or subject entity. Then, we obtain the mask embedding $\mathbf{e}_{[mask]}^{(l)}$ through Eq. (2), and feed it into the multi-layer perceptron (MLP) decoder to maximize the probability of the missing entity $o$ and $s$ through the cross-entropy loss function:
$p\left(\mathbf{e}\right)=\text{Softmax}(\mathbf{e}_{[mask]}^{(l)}\mathbf{w}+\mathbf{b}),$ (5)
$\mathcal{L}=-\sum_{(s,r,o,t)\in\mathcal{G}}\log p({o_{t}})+\log p({s_{t}}).$ (6)
4.2.2 Virtual Knowledge Indicators
We design three knowledge indicators to link graph signals and input prompt text, namely [SUB], [REL] and [OBJ], correspond to the virtual tokens of the head entities, relations and tail entities in the subgraph, respectively. We then try to incorporate structural and temporal information from the subgraph into the indicator representations. Specifically, we use the $\operatorname{Local}$ operator to get the structure representations $\mathbf{h}^{s}$ of entities and relationships in the subgraph, respectively:
$\mathbf{h}_{\operatorname{[SUB]}}^{s}=\operatorname{Local}(\mathbf{e}_{\operatorname{[SUB]}}),$ (7)
here $\mathbf{e}_{\operatorname{[SUB]}}$ represents pre-trained T-GNN embeddings of all subject entities in the subgraph, $\operatorname{Local}$ indicates the max or mean pooling operator. Besides, we leverage the time embeddings to enhance the indicator representations with temporal information.
$\mathbf{h}_{\operatorname{[SUB]}}^{st}=\mathbf{h}_{\operatorname{[SUB]}}^{s}+\mathbf{t}_{min}+\mathbf{t}_{max},$ (8)
where $\mathbf{t}_{min}$ and $\mathbf{t}_{max}$ denote the embeddings for the minimu and maximum values of time in the subgraph, respectively. The intuition follows BERT that use position embeddings for tokens Devlin et al. (2019). Here, time embeddings can be seen as entity positions in the time dimension. The relation and object indicators are the same as subject. At last, we employ a simple linear layer $\mathbf{W}_{p}$ to project them into the textual representation space of the LLM. The final input prompt sequence $\mathcal{S}=\mathcal{V}:\mathcal{I}:\mathcal{Q}:\mathcal{A}$, $\mathcal{V}$ represent virtual indicator tokens. Details of the instruction template can be found in the Appendix C. The optimization objective of the LLM $\mathcal{M}$ can be formulated as:
$\mathcal{A}=\arg\max_{\mathcal{A}}P_{\mathcal{M}}\left(\mathcal{A}\mid\mathcal{V},\mathcal{I},\mathcal{Q},\mathcal{A}\right).$ (9)
Model Hits@1 Hits@10
Overall Question Type Answer Type Overall Question Type Answer Type
Complex Simple Entity Time Complex Simple Entity Time
EmbedKGQA 0.288 0.286 0.290 0.411 0.057 0.672 0.632 0.725 0.850 0.341
EaE 0.288 0.257 0.329 0.318 0.231 0.678 0.623 0.753 0.668 0.698
CronKGQA 0.647 0.392 0.987 0.699 0.549 0.884 0.802 0.990 0.898 0.857
EntityQR 0.745 0.562 0.990 0.831 0.585 0.944 0.906 0.993 0.962 0.910
TMA 0.784 0.632 0.987 0.792 0.743 0.943 0.904 0.995 0.947 0.936
TSQR 0.831 0.713 0.987 0.829 0.836 0.980 0.968 0.997 0.981 0.978
TempoQR 0.918 0.864 0.990 0.926 0.903 0.978 0.967 0.993 0.980 0.974
BERT w/o tkg 0.071 0.086 0.052 0.077 0.06 0.213 0.205 0.225 0.192 0.253
RoBERTa w/o tkg 0.07 0.086 0.05 0.082 0.048 0.202 0.192 0.215 0.186 0.231
ChatGPT w/o tkg 0.151 0.144 0.160 0.134 0.182 0.308 0.308 0.307 0.257 0.402
BERT w/ tkg 0.243 0.239 0.249 0.277 0.179 0.620 0.598 0.649 0.628 0.604
RoBERTa w/ tkg 0.225 0.217 0.237 0.251 0.177 0.585 0.542 0.644 0.583 0.591
ChatGPT w/ tkg 0.754 0.579 0.987 0.689 0.873 0.852 0.746 0.992 0.808 0.933
GenTKGQA 0.978 0.962 1.000 0.964 0.999 0.981 0.968 1.000 0.971 0.999
Table 1: Performance comparison of different models on CronQuestions. The best and second best results are marked in bold and underlined, respectively. w/o tkg indicates that LMs answer the questions directly without using TKG information, and w/ tkg indicates that LMs answer the questions with TKG background knowledge.
5 Experiments
We design experiments to answer the following questions:
Q1.How does GenTKGQA perform on the TKG question answering task? (Section 5.2)
Q2.How do the two stages contribute to the model performance respectively? (Section 5.3)
Q3.How does GenTKGQA perform under changes in hyper-parameters? (Section 5.4)
Q4.How does GenTKGQA outperform ChatGPT in answering complex temporal questions? (Section 5.5)
5.1 Datasets, Metrics and Baselines
CronQuestions Saxena et al. (2021) is a temporal QA dataset based on the Wikidata TKG Lacroix et al. (2020), comprising 125k entities, 203 relations, 1.7k timestamps and 328k facts. It contains 410K unique question-answer pairs, including annotated entities and timestamps, with 350k for training and 30k for validation and testing. The dataset can be categorized into simple reasoning (Simple Entity and Simple Time) and complex reasoning (Before/After, First/Last, and Time Join) based on temporal constraints. Following previous studies, we use two popular evaluation metrics, Hits@1 and Hits@10. More information about datasets and metrics can be found in Appendix A.
We compare four types of baselines: 1) KG embedding-based models including EaE Feng et al. (2020b) and EmbedKGQA Saxena et al. (2020); 2) TKG embedding-based models including CronKGQA Saxena et al. (2021), EntityQR Mavromatis et al. (2022), TMA Liu et al. (2023a), TSQA Shang et al. (2022) and TempoQA Mavromatis et al. (2022); and 3) Language models, including BERT Devlin et al. (2019), RoBERTa Liu et al. (2019), and ChatGPT OpenAI (2023). The implementation details of the baselines and our model are described in Appendix B.
5.2 Main Results
Table 1 reports the performance of all methods on the CronQuestions dataset for various question types. We can observe that GenTKGQA consistently outperforms the baselines in terms of "Overall" performance, and achieves significant improvements of 11.3% in the "Complex" question type and 10.6% in the "Time" answer type on the Hits@1 metric over the second best method. Especially, our model achieves 100% for the "Simple" question type. The possible reason is that "Simple" questions usually involve single facts, GenTKGQA can easily retrieve the relevant facts containing the answer and infer the correct answer through instruction tuning technique.
Furthermore, compared to KG embedding methods, temporal KG embedding methods show significant results on various metrics, thanks to the fact that the temporal information of the TKG is taken into account in the question representation. This is also why KG embedding methods are particularly ineffective for the "Time" answer type. However, most TKG embedding methods treat the QA task as a link prediction, which works for the "Simple" question type containing a single fact compared to the "Complex".
We find that PLMs (BERT, RoBERTa) and LLMs (ChatGPT) have the lowest performance on the TKGQA task without TKG information. This suggests that language models (LM), whether encoded or generated, with a large or small number of parameters, have difficulty answering temporal questions without any relevant context. w/ tkg indicates that LMs use entity/time embeddings or relevant facts from the TKG. Obviously, LMs w/ tkg have significantly better performance, which suggests that the language models have some degree of temporal reasoning capability when relevant TKG information is provided, validating the importance of the subgraph retrieval phase. It is worth noting that ChatGPT w/ tkg still performs weakly in reasoning about complex problem types when providing the facts retrieved in the first stage by GenTKGQA, while our model achieves the best results. This demonstrates the effectiveness of interacting GNN and LM representations in a non-shallow way in dealing with complex temporal problems. The above findings demonstrate the adequacy of our two motivations for solving complex temporal problems with LLMs. A possible reason for the poor improvement of our model’s Hits@10 metric for the "Entity" type is that the generative LMs generate irrelevant responses when asked to generate multiple answers. Last but not least, GenTKGQA, as a generative QA model, achieves better results than traditional extractive QA methods.
Model Hits@1
Overall Question Type Answer Type
Complex Simple Entity Time
GenTKGQA 0.978 0.962 1.000 0.964 0.999
w/o SR 0.119 0.140 0.090 0.127 0.103
w/o SR inference 0.475 0.381 0.601 0.294 0.812
w/ SR random 0.766 0.613 0.970 0.661 0.961
w/o T-GNN 0.935 0.914 0.965 0.920 0.963
w/o VKI 0.843 0.824 0.870 0.831 0.867
Table 2: Ablation study results on CronQuestions.
5.3 Ablation Study
As shown in Table 2, to verify each module’s importance, we conduct ablation experiments on the CronQuestions dataset.
w/o SR means that we directly perform problem inference without using relevant subgraph information in both the model training and inference phases, while w/o SR inference means that we do not provide subgraphs only at the inference. We can observe a sharp decrease in model effectiveness due to the lack of use of subgraph information, which is consistent with the other LMs (Section 5.2). This result shows that the current LMs do not have the reasoning capability to deal directly with temporal reasoning problems, validating the importance of the subgraph retrieval module. The w/o SR inference result indicates that GenTKGQA remembers part of the structured knowledge during the training phase and improves the temporal inference performance without providing subgraph information. w/ SR random denotes the random selection of relevant facts involving entities in the question, and the drop in results proves the validity of our first-stage approach.
w/o T-GNN indicates that we directly use temporal embeddings to represent entities and relations of subgraphs. The slight decrease in the results indicates that the T-GNN is able to perceive the structural information of the TKG. w/o VKI means that we try to remove the virtual knowledge indicators from the input prompt. Model performance degradation shows that indicators can bridge the gap between distinct representations.
Figure 3: Parameter sensitivity on our GenTKGQA.
Type Question/ Response
Retrieved Graph ChatGPT w/ tkg GenTKGQA
Simple dean in 1997 was the person?
Entity [Xavier Darcos, position held, dean, 1995, 1998] [ Katarzyna Olbrycht, José Miguel Pérez García, [ Xavier Darcos, Zinaida Belykh,
[Zinaida Belykh, position held, dean, 1988, 1998] Jiří Zlatuška, Xavier Darcos, José Miguel Pérez García, Jiří Zlatuška,
[José Miguel Pérez García, position held, dean, 1990, 1998] Zinaida Belykh, Andrei Fursenko, Katarzyna Olbrycht, Catalina Enseñat Enseñat,
[Jiří Zlatuška, position held, dean, 1994, 1998] Anatoly Torkunov, Alexander Konovalov, Catalina Enseñat Enseñat, Marcel Berger,
[Katarzyna Olbrycht, position held, dean, 1981, 1998] Anatoly Vichnyakov, Anatoly Vishnevsky] Miguel Beltrán Lloris, Miklós Réthelyi]
Simple When Daniele Amerini played in Modena F.C.?
Time [Daniele Amerini, member of sports team, Modena F.C., 2005, 2006] [2005, 2006, [2005, 2006,
[Daniele Amerini, member of sports team, Modena F.C., 2008, 2009] 2008, 2009] 2008, 2009]
Before/ What was the first General Inspector
After of the Bundeswehr after 2004?
[Wolfgang Schneiderhan, position held, Bundeswehr Chief, 2002, 2009] [Volker Wieker] [Wolfgang Schneiderhan]
[Volker Wieker, position held, Bundeswehr Chief, 2010, 2018]
First/ When Mark Burke was playing his final game?
Last [Mark Burke, member of sports team, Luton Town F.C., 1994, 1994]
[Mark Burke, member of sports team, Port Vale F.C., 1994, 1995] [1994] [1995]
[Mark Burke, member of sports team, Wanderers F.C., 1991, 1994]
[Mark Burke, member of sports team, Darlington F.C., 1990, 1990]
Time Who worked with Christiane Hoffmann
Join during his employment with FAZ? [Gero von Randow, Frank Schirrmacher,
[Christiane Hoffmann, employer, FAZ, 1994, 2010] [Christiane Hoffmann] Oliver Gehrs, Peter Kiefer,
[Gero von Randow, employer, FAZ, 2001, 2003] Wolfgang Bresser, Ulrich Kuehl
[Frank Schirrmacher, employer, FAZ, 1994, 2014] Henning Franke, Stephan Löwenstein]
Table 3: Comparison of responses to five different question types between our GenTKGQA and ChatGPT w/ tkg. Marked in blue is the correct answer.
5.4 Sensitivity Analysis
Impact of training data size.
We explore the impact of different training data sizes to reason about complex temporal questions. As shown in Figure 3(a), by comparing the Hits@1 metric of several methods for the complex question type, our method consistently outperforms others as the training data expands. In particular, at 20% of the training data, our model outperforms the second best model by 32.3%, demonstrating that our model has strong inference ability in the case of few-shot samples due to its intrinsic knowledge.
Impact of the number of relevant facts.
We report the performance changes on the CronQuestions dataset by varying the number of retrieved facts $n$ in Figure 3(b). It can be seen that the model performs poorly with a small number of relevant facts ($n$=3), and there is a slight drop in performance at $n$=8. Fewer facts do not provide sufficient context knowledge, while more facts may introduce noise. Taking this into consideration, we set the hyper-parameter $n$ to 5.
5.5 Qualitative Results
We provide specific examples for each question type to compare the answer results of ChatGPT and ChatGPT w/ tkg. Table 3 includes the graphs retrieved by our method, along with the answer results for five different question types.
When providing relevant facts retrieved by GenTKGQA as background knowledge, ChatGPT performs competitively in the simple question type, correctly answering questions with entity or time as the answer. However, it has difficulty answering complex types of questions. For example, in the "Before/After" and "First/Last" question types, ChatGPT struggles to understand the temporal order of the relevant facts in the time dimension and gives incorrect answers. Besides, it does not seem to understand the dependencies among related facts. In the case of the "Time Join" question type, which asks for other entities who worked with Christiane Hoffmann at a certain time, ChatGPT wrongly considers the entity Christiane Hoffmann itself that appears in the relevant facts as the correct answer. On the contrary, GenTKGQA performs well in both simple and complex question types, thanks to the fact that our proposed approach allows the LLM to perceive the structural and temporal information of the retrieved subgraphs. However, similar to other generative LLMs, GenTKGQA randomly generates some irrelevant answers when generating multiple answers, e.g., in the "Simple Entity" and "Time Join" question types.
6 Conclusion
We propose a novel generative framework, GenTKGQA, which guides the LLM in a two-stage manner to handle temporal question answering on TKG. Specifically, at the subgraph retrieval phase, we adopt a divide-and-conquer strategy to exploit the LLM’s intrinsic knowledge to mine the temporal constraints and structural links in the temporal questions to reduce the search space of the subgraphs in both time-space dimensions. We employ the in-context learning approach to complete subgraph retrieval for the entire dataset with a small number of samples. In order to improve the inference performance of the LLM on complex types of problems, at the answer generation phase, we present the instruction tuning technique to make the open-source LLM truly understand the temporal order and structural dependencies among retrieved facts. Most critically, we design novel knowledge indicators to establish a bridge between subgraph neural information and text representations. Experimental results show that our framework can effectively utilize the LLM to solve the complex question type of TKGQA task and validate the adequacy of our motivation.
Limitations
Although the complex temporal question in the CronQuestions dataset involves multiple facts, the inter-entity connection in each fact is single-hop, so the hyper-parameter $k$ of our model is set to 1 to achieve the best results. In fact, the vast majority of current TKGQA datasets involve facts that are single-hop. So, in the future, we will explore more datasets to solve inference for multi-hop complex temporal problems over TKG.
Appendix A Dataset Statistics and Metrics
We use the CronQuestions dataset in our experiments. Dataset statistics and TKG information are described in Table 4 and 5, respectively.
[…]
Appendix B Baselines and Implementation Details
We use the OpenAI-API [^1] https://platform.openai.com/docs/api-reference (gpt-3.5-turbo-0613222https://platform.openai.com/docs/models/gpt-3-5-turbo) for all ChatGPT-related experiments, including subsequent ChatGPT baselines.
In the subgraph retrieval phase, we use ChatGPT OpenAI (2023) to mine temporal constraints and structural links between entities and add 5 samples to the in-context learning prompt templates, which are presented as Table 6 and 7. We set $k$=1 for the top-$k$ relations. In the answer generation phase, following Lacroix et al. (2020), we select the dimension of entity/relation/time embeddings to 512. For T-GNN, the layer $l$ is set to 1, the linear transformations $\mathbf{W}_{q}$, $\mathbf{W}_{k}$ and $\mathbf{W}_{m}$ are 512×512, and the $\mathbf{m}$ and $\mathbf{b}$ of the MLP layer are $|\mathcal{E}|$. We use the open-source Llama 2-7B Touvron et al. (2023) for instruction tuning and select up to $n$=5 relevant facts as additional knowledge. The linear projection layer $\mathbf{W}_{p}$ is 512×4096. We fine-tune Llama 2-7B using LoRA Hu et al. (2022) with rank 64. The number of epochs is set to 3 and the learning rate is 3e-4. We use the AdamW optimizer Loshchilov and Hutter (2019) with a fixed batch size of 8. We conduct all the experiments with NVIDIA A100 GPUs, and the results of each experiment are averaged over three runs. We will release the source code upon acceptance.
We compare our model with the following baselines:
EmbedKGQA Saxena et al. (2020): Timestamps are ignored during pre-training and random time embeddings are used during the QA task.
EaE Feng et al. (2020b): In the experiment, we follow use TKG embeddings to enhance the question representation, and then predict the answer probabilities via dot-product.
CronKGQA Saxena et al. (2021): CronKGQA is the TKGQA embedding-based method that first uses a LM model to get question embeddings and then utilize a TKG embedding-based scoring function for answer prediction.
EntityQR and TempoQR Mavromatis et al. (2022): Based on EaE, EntityQR utilizes a TKG embedding-based scoring function for answer prediction. TempoQR utilizes a TKG embedding-based scoring function for answer prediction and fuse additional temporal information.
TMA Liu et al. (2023a): TMA improves QA performance through enhanced fact retrieval and adaptive fusion network.
TSQR Chen et al. (2021b): TSQA presents a contrastive learning module that improves sensitivity to time relation words.
BERT and RoBERTa Devlin et al. (2019); Liu et al. (2019): For w/o tkg, following CronKGQA Saxena et al. (2021), we add a prediction head on top of the [CLS] token of the final layer, and then do a softmax over it to predict the answer probabilities. For w/ tkg, following TempoQR Mavromatis et al. (2022), we generate their LM-based question embedding and concatenate it with the annotated entity and time embeddings, followed by a learnable projection. The resulted embedding is scored against all entities and timestamps via dot-product.
ChatGPT OpenAI (2023): To ensure that the output format meets the expected requirements, we use the in-context learning approach to motivate ChatGPT to answer the questions and provide 5 examples in the prompt template. The specific templates are presented in Table 9 of Appendix C. w/o tkg and w/ tkg differ in whether or not question-relevant facts are provided in the input prompts, which are retrieved in the first stage by our proposed GenTKGQA framework.
Relation Ranking Prompt
l will give you a list of words.
Find the $\{k\}$ words from the list that are most semantically related to the given sentence.
If there are no semantically related words, pick out any $\{k\}$ words.
Examples)
Sentence A: When was the first time Martin Taylor played for The Hatters?
Words List: [‘member of sports team’, ‘position held’, ‘award received’, ‘spouse’, ‘employer’]
Top {$k$} Answers: [‘member of sports team’]
Sentence E: Which was awarded to Daniel Walther in 1980?
Words List: [‘member of sports team’, ‘position held’, ‘award received’, ‘spouse’, ‘employer’]
Top {$k$} Answers: [‘award received’]
Now let’s find the top {$k$} words.
Sentence: {$sentence$}
Words List: {$relation\_list$}
Top {$k$} Answer:
Table 6: Relation Ranking Prompt. This prompt is used to extract structural links between entities in the question.
Time Mining Prompt
I will give you a natural language question with a temporal constraint.
Answer the temporal constraint involved in the question based on the knowledge context and the question type.
Answer only in "before", "after", "between and" format.
Examples)
Question A: Who held Governor of Connecticut position after Lowell P. Weicker?
Knowledge Context: [‘Lowell P. Weicker’, ‘position held’, ‘Governor of Connecticut’, ‘1991’, ‘1995’]
Question Type: after
Response: after 1995
Question E: Who’s the player who played in AC Reggiana with Daniele Magliocchetti?
Knowledge Context: [‘Daniele Magliocchetti’, ‘member of sports team’, ‘A.C. Reggiana’, ‘2012’, ‘2014’]
Question Type: time_join
Response: between 2012 and 2014
Next, let’s answer the time constraints involved in the following question.
Question: {$question$}
Knowledge Context: {$context$}
Question Type: {$type$}
Response:
Table 7: Time Mining Prompt. This prompt is used to find the time constraints involved in the complex question.
Instruction Tuning Template
Below is an instruction that describes a task, paired with an input that provides further context.
Write a response that appropriately completes the request.
Instruction:
Answer the questions based on evidence.
Each evidence is in the form of [head, relation, tail, start_time, end_time]
and it means ‘head relation is tail between start_time and end_time’.
You must list the 10 most relevant answers.
Input:
Question: {$question$}
Evidence set: {$evidence\_set$}
Response:{$answer$}
Table 8: This is the template for instruction tuning.
ChatGPT w/ tkg
Answer the questions based on evidence.
Each evidence is in the form of [head, relation, tail, start_time, end_time]
and it means ‘head relation is tail between start_time and end_time’.
[…]
Examples)
Question A: Who was the Member of the House of Representatives in 1990?
Evidence set: [[‘Simon Crean’, ‘position held’, ‘Member of the House of Representatives’, ‘1990’, ‘2013’],…]
[…]
Question E: With whom did Steve Haslam play on the Sheffield Wednesday F.C.?
Evidence set: [[‘Ola Tidman’, ‘member of sports team’, ‘Sheffield Wednesday F.C.’, ‘2003’, ‘2005’], …]
[…]
Now let’s answer the Question based on the Evidence set.
[…]
Question: {$question$}
Evidence set: {$evidence\_set$}
Answer:
ChatGPT w/o tkg
Answer the questions directly.
[…]
Examples)
Question A: Who was the Member of the House of Representatives in 1990?
[…]
Question E: With whom did Steve Haslam play on the Sheffield Wednesday F.C.?
[…]
Question: {$question$}
Answer:
Table 9: ChatGPT Baseline Prompt.
Appendix C Prompt Template
The prompts for relation ranking and time mining can be found in Table 6 and Table 7, respectively. The template used for instruction tuning is shown in Table 8. The ChatGPT baseline prompt is presented in Table 9.